### v1

# Detecting Document-level Paraphrased Machine Generated Content: Mimicking Human Writing Style and Involving Discourse Features

Abstract
The availability of high-quality APIs for Large Language Models (LLMs) has facilitated the widespread creation of Machine-Generated Content (MGC), posing challenges such as academic plagiarism and the spread of misinformation. Existing MGC detectors often focus solely on surface-level information, overlooking implicit and structural features. This makes them susceptible to deception by surface-level sentence patterns, particularly for longer texts and in texts that have been subsequently paraphrased.
To overcome these challenges, we introduce novel methodologies and datasets. Besides the publicly available dataset Plagbench, we developed the paraphrased Long-Form Question and Answer (paraLFQA) and paraphrased Writing Prompts (paraWP) datasets using GPT and DIPPER, a discourse paraphrasing tool, by extending artifacts from their original versions. To address the challenge of detecting highly similar paraphrased texts, we propose MhBART, an encoder-decoder model designed to emulate human writing style while incorporating a novel difference score mechanism. This model outperforms strong classifier baselines and identifies deceptive sentence patterns. To better capture the structure of longer texts at document level, we propose DTransformer, a model that integrates discourse analysis through PDTB preprocessing to encode structural features. It results in substantial performance gains across both datasets – 15.5% absolute improvement on paraLFQA, 4% absolute improvement on paraWP, and 1.5% absolute improvemene on M4 compared to SOTA approaches.
Detecting Document-level Paraphrased Machine Generated Content: Mimicking Human Writing Style and Involving Discourse Features
Yupei Li1, Manuel Milling2, Lucia Specia1, Björn W. Schuller1,2, 1GLAM team, Imperial College London, London, UK, 2CHI – Chair of Health Informatics, Technical University of Munich, Munich, Germany,
1 Introduction
As LLMs such as GPT-4 OpenAI (2023) and Llama3 Touvron et al. (2023) develop rapidly and become widely utilized by the general public, Machine-Generated Content (MGC) has become increasingly abundant. In particular, people often rely on these models to revise long texts, such as essays, which results in paraphrased MGC. One example is shown in Figure 1, where the original human-written text is highlighted in red on the left side, and the MGC in green on the right side, with differences shaded. It can be argued that content features such as fluency, readability, and clarity all improve after revision, and that the usage of LLM-usage can be hard to detect manually.
Figure 1: Comparison of human-written text (left) with its corresponding GPT-revised text (right).
There are well known risks associated with the misuse of MGC. Crothers et al. (2023) conducted a survey on the applications of MGC, reporting that it may contribute to academic misconduct, the spread of fake political news, and an increase in social media spam. Our goal in this paper is the effective detection of document-level paraphrased MGC, which is rewritten by LLMs from an original human written text.
Automatically distinguishing MGC from Human-Produced Content (HPC) is, however, very challenging, given that MGC tends to resemble HPC, often containing at least chunks that are reproduced from human produced training data. Currently, methods to detect MGC involve a combination of processes, namely a Natural Language Understanding (NLU) module that extracts linguistic and contextual features, followed by a classification module which fits this to binary labels using algorithms such as Logistic Regression (LR) Cox (1958) for decision-making.
However, these suffer from various limitations. First, current NLU models built using Transformers Vaswani et al. (2017) such as Roberta Liu et al. (2019) demonstrate good performance in analysing short texts (such as sentences) but less so when processing document-level data, as shown in Fu (2024). They tend to struggle with retaining information if the texts are long. Even LLMs are limited in the length of the input texts that they can process effectively. Document-level texts exhibit more complex structures’ relationships. Secondly, the degree of similarity between MGC and HPC is higher especially when LLMs are used to paraphrase HPC rather than generating it from scratch. This makes distinguishing HPC and MGC challenging, as differences in linguistic content are less distinctive. Previous work focuses on identifying similarities at surface level linguistic features, ignoring the importance of other implicit features such as writing style or structural features, which we show can contribute to the effectiveness of a detection model. There is indeed a difference in writing style in Figure 1: the left side is more poetic, while the right side is a plainer narrative, though the distinction may not be obvious to the human reader. These limitations lead to underperforming detectors in work to date.
We propose two novel approaches to automatically detect MGC which address the limitations described above. The aim is to better classify the texts by identifying structural features, and implicit features, such as discourse relations or writing styles. Specifically, we introduce two models: MhBART, which does detection through writing style by leveraging the capabilities of BART Lewis et al. (2019), and DTransformer, which incorporates discourse features to improve detection accuracy.
As data to build and evaluate our models, we use publicly available datasets containing document-level paraphrased MGC and their HPC counterparts, as well as two new datasets that we created. We show that our proposed models consistently outperform baseline and SOTA approaches, both research and commercial tools, for long-form and machine-paraphrased texts. Our contributions can be summarized as follows:
A. We create paraphrased Long-Form Question and Answer (paraLFQA) and paraphrased Writing Prompts (paraWP) – two parallel HPC-MGC corpora for model evaluation and further research in Section 3
B. We introduce MhBART – an encoder-decoder model designed to learn human writing style, which outperforms strong baselines for paraphrased MGC detection in Section 4. Additionally, we identified sentence patterns that can deceive current detectors.
C. We introduce DTransformer – a discourse-aware Transformer-based model to incorporate discourse features in detection, which leads to better performance for long form texts in Section 5.
2 Related work
2.1 MGC Detection Approaches
Given the current prominence of this topic, current work includes contributions from both academia and industry.
In the academic literature, two types of models are commonly used: statistical methods, also known as feature-based methods, and neural network methods. Herbold et al. (2023) identified seven linguistic features and criteria for distinguishing MGC from HPC, and they concluded that MGC tends to exhibit fewer discourse and epistemic words but displays greater lexical diversity and nominalizations. Choudhary and Arora (2021) and GLTR (Generative Language Test Room) Gehrmann et al. (2019) also made decision with linguistic features.
However, feature-based methods need manual work to design the features. There have been studies comparing neural network methods and statistical methods Hayawi et al. (2023), demonstrating the superior efficiency of Transformer-based models Vaswani et al. (2017), such as Roberta Large OpenAI Detector Solaiman et al. (2019), the first fake news detector Grover Zellers et al. (2019), a zero-shot detection method DetectGpt Mitchell et al. (2023), watermarking based detection Fu et al. (2024). The Roberta Large Detector has been fined-tuned with 1.5B output of GPT2. The detection is treated as a binary classification. However, it was trained on texts produced with GPT-2, which may retain biases.
Additionally, Fu et al. (2024) have introduced semantic-aware watermarking algorithms, while Jiang et al. (2023) employ perturbation strategies to evade detection methods based on watermarking. Besides, Hu et al. (2023) introduced another adversarial learning approach named Radar, noted for its robust transferability. Additionally, Chaka (2023) directly utilizes LLMs Brown et al. (2020) for MGC detection, but concludes that these models are currently inadequate. Conversely, Chen et al. (2023) combine two LLMs, BERT Devlin et al. (2019), and T5 Raffel et al. (2020), to improve detection efficacy. Moreover, recent advancements include the M4 model for multi-modality MGC detection Wang et al. (2023) and DeBERTa He et al. (2020) for monitoring mainstream news websites Hanley and Durumeric (2024).
Moreover, numerous widely utilized commercial tools commercial tools include GPTZero  [^1] https://gptzero.me/, Copyleaks [^2] https://copyleaks.com/ai-content-detector, Content at Scale  [^3] https://contentatscale.ai/ai-content-detector/, Originality  [^4] https://originality.ai/, Winston AI  [^5] https://gowinston.ai/, and Hive  [^6] https://hivemoderation.com/ai-generated-content-detection. However, these tools utilize models that are not publicly accessible, which makes understanding these tools challenging.
As discussed above, the open challenges that have not covered in previous work are as follows: Combining implicit features with linguistic features, and detecting document-level paraphrased machine-generated texts.
2.2 Detection of Paraphrased-MGC
MGCs learn from human texts which have similar features as HPCs, so MGCs and HPCs become more similar, especially for machine-paraphrased MGC datasets. Krishna et al. Krishna et al. (2023) developed an attack method targeting recent detectors by employing paraphrased sentences in MGC, resulting in a significant decrease in detector accuracy. This approach utilized paraphrased sentence patterns introduced by an external model known as DIPPER Krishna et al. (2023), highlighting the challenges associated with detecting discourse-paraphrased sentences.
Efforts have been made to address the paraphrased MGC detection issue. Wahle et al. (2022) and Foltỳnek et al. (2020) proposed using pre-trained word embedding models for text representation, combined with classifiers, to predict paraphrased plagiarism in writing. Animesh et al.  Nighojkar and Licato (2021) introduced an adversarial method to detect paraphrasing by leveraging the inferential properties of sentencesOther research has focused on developing corpus-based methods for detecting paraphrasing Vrbanec and Meštrović (2020). However, paraphrasing involves more than just substituting words and should incorporate additional detectable features. Researchers have manually crafted features, as discussed by Shahmohammadi et al.  Shahmohammadi et al. (2021), which align with syntactic similarity, and other factors.
The aforementioned methods primarily focus on content similarity detection. However, these sentences having formats of discourse paraphrasing are hard to be addressed, which involve changing the structure of sentences and the organization of paragraphs.
2.3 Detection of Document-level MGC
Long texts have more complex language organization compared to shorter texts, so comprehending long texts is more challenging than shorter ones. Current Transformer-based models can fail when processing longer sequences Wolf et al. (2020) and there remains a risk of information loss due to the constrained memory and computational costs associated with self-attention operations Xie et al. (2023).
To address the aforementioned issue, hierarchical models have been proposed that handle sentence-level information before passing on these representations for document-level processing. Zhang et al. (2019) proposed the HIBERT model for document summarization. Moreover, Xie et al. (2023) divided long sequences into different blocks, while SPECTER Cohan et al. (2020) was proposed to embed research papers by utilizing only the metadata. Longformer Beltagy et al. (2020) introduced a sparse attention mechanism that adapts to sequence length, and Grail et al. (2021) proposed to globalize sequences processed by BERT using a bidirectional GRU network to propagate information. Each of these models aim to address memory limitations and computational complexity by selecting representative embeddings. However, there are no dominating models utilizing discourse features.
The discourse features could be represented by two annotations: Rhetorical Structure Theory Discourse Treebank (RST-DT) Carlson et al. (2006), a dataset that captures discourse relations within text spans, represented in a tree structure and the Penn Discourse Treebank (PDTB) Prasad et al. (2008), offering a straightforward representation of relations, specifying two arguments along with their relations Webber et al. (2019) (as depicted in Figure 2). Annotated by human experts, PDTB’s 3.0 version includes over 100 types of relations, demonstrating exceptional applications when combined with Bert models Kim et al. (2020). These annotated labels are detailed about the role of each sentences as well.
Figure 2: PDTB annotation example Lin et al. (2010): It shows the implicit relations, ACCORDINGLY, between the given two sentence.
3 Datasets
Currently, limited public datasets are available that focus on document-level paraphrased texts for MGC detection. For example, the recently published adversarial perturbations dataset introduced revised MGC texts with word replacements Huang et al. (2024), but the dataset does not include meaningful differences in writing style between the two versions. We present the Plagbench Lee et al. (2024) and Multi-domain, and Multi-lingual Black-Box Machine-Generated Text Detection (M4) dataset Wang et al. (2024) as publicly available datasets, and also introduce two new dataset that we created based on a portion of well-known datasets used in other NLP tasks: LFQA Fan et al. (2019), which was also used in the DIPPER dataset Krishna et al. (2023) and WritingPrompts (WP) Fan et al. (2018).
The Plagbench dataset Lee et al. (2024) is generated by GPT specifically for the task of plagiarism detection. We select the paraphrasing plagiarism subset, which consists of 135 data pairs.
The M4 dataset contains a diverse array of human-written sources, including document-level texts from WikiHow, and others. It also includes MGC versions produced by a wide range of generators, such as BLOOMz Team (2022) and ChatGPT. These machine-generated texts often differ in meaning from the original sources because the prompts used are not direct paraphrases but involve expansions, such as the prompt “Write a peer review based on the article." It has been extensively utilized as a benchmark in MGC detection research. The dataset has a total of 133,551 text samples.
For the paraLFQA dataset, the original answers from LFQA are considered as HPC. MGC was produced using the DIPPER model from the HPCs, resulting in discourse-paraphrased-style sentences, which we present as a new dataset. In accordance with the methodology outlined in the original paper, we present T5 as the tokenizer. We selected a lexicon code of 40 and a re-order code of 100. These values represent that 40% of the lexicon and 100% of the order have been modified. These hyperparameters were chosen to emphasize changes in the organization of the language within the dataset. The dataset has around 2,000 MGC-HPC paragraph pairs with around 250-350 words in each paragraph. This serves as a control dataset where the semantic features are similar between the paraphrased and original content, but the writing style is different.
For the paraWP dataset, human-written stories from the provided prompts are considered as HPC. The MGC paraphrases are produced by GPT-3.5-turbo with the following prompt: “Summarize the text. Expand the idea into new text. Keep the original narrative, such as he, she, I, or their names. Return the same number of words as the input texts." In this way, the data has one-to-one matched text pairs with HPC and MGC, which are regarded as paraphrased-style versions. We have created 166,247 documents both for HPC and MGC in the training set, and 12,800 and 12,316 documents for the validation and test sets, respectively. We used BPE tokenizer to process this data and the vocabulary size of the BPE modelling is 8,016. We have also investigated linguistic features of the dataset in Appendix A.
Both paraLFQA and paraWP are downloadable [^7] https://drive.google.com/file/d/1fvsWwHKplf0-n6PnwbxIRmR6jgu62nRi/view?usp=sharing and are licensed under a Creative Commons Attribution 4.0 International License.
4 MhBART: Mimicking human writing style Bidirectional and Auto-Regressive Transformers
This section introduces our first method especially for the first challenge: detecting MGC with high similarity to the original text. MhBART focuses on learning to mimic human writing style, aiming to detect MGC by identifying deviations from natural language patterns typically found in human-authored content. The method reverses the process of generating MGC in LLMs like GPT, i. e., our model is learning to generate HPC from MGC, such that we can implicitly grasp the writing style of humans.
The motivation behind this approach lies in the capability of language models to comprehend language, enabling them to learn how humans organize words and their overall written style, which constitutes the implicit features of the text. By training the model to transfer from MGC to HPC, the expectation is that it could capture and replicate human-like writing patterns and style effectively. We selected BART Lewis et al. (2019) as the base pre-trained model due to its ability to handle long texts.
The overall model architecture is illustrated in Figure 3. It learns from pairs of MGC and HPC samples that have nearly identical content and ideas but differ in their organization or style. For training, the input to the MhBART consists of a document from this MGC-HPC paired set, and the target (or supervision) is the corresponding HPC. This setup is reminiscent of a translation model, where the source-side sentences are MGC and the target-side sentences are HPC.
During inference and classification, the trained model evaluates the difference between inputs and outputs using various string similarity metrics. Specifically, we employ BLEU Papineni et al. (2002), which assesses similarities in form at the n-gram level, and cosine similarity score, which measures similarity of representations in a shared embedding space and assesses more features based on what the embedding represents. We calculate the cosine similarity score based on the representations of sentences obtained from BERT. We also investigate the cosine similarity score calculated based on universial authorship representation (UAR) Soto et al. (2021) but it shows quite negative results (details discussed in Appendix A). We deem this phenomenon is attributed to the limitations of UAR in effectively capturing the entire writing style and its challenges in fine-tuning for our specific task, so we did not utilize that. We use the four components of the BLEU score and the cosine similarity score as features for a logistic regression (LR) classification model. This LR model learns the boundary conditions necessary to classify texts as either HPC or MGC.
The hypothesis is that BART, when trained to generate texts that adhere to human-written style, will produce outputs that differ significantly from input texts if they are MGC as the model is trained to learn HPC writing style. Therefore, if the model receives HPC input, it would not change the writing significantly.
Figure 3: MhBART approach: Learning to translate from MGC to HPC. The model emulates HPC writing style by training with BART with the input being the corresponding MGC. An intermediate output, which captures properties indicative of human writing style, is generated by the trained BART model and subsequently compared with the input text. A substantial difference score suggests the text is MGC, whereas a low difference score supports classification as HPC.
5 DTransformer
This section introduces our second method, specifically designed to address the second challenge: detecting document-level MGC texts paraphrased from HPC. The DTransformer model focuses on integrating structural features with semantic features, aiming to detect MGC by identifying the ways humans and machines organise text blocks, such as paragraphs. This model is inspired by the discourse information extraction method with PDTB annotation and the effects of structural features discussed in the Multi-Method Qualitative Text Framework Alejandro and Zhao (2023). The model is depicted in Figure 4, with detailed sub-modules discussed in the following paragraphs.
Figure 4: DTransformer model: it incorporates both structural and semantic features. It first splits the paragraphs and employs a hierarchical model to capture document semantic features. Then, using a cross-attention mechanism, it analyzes discourse features for improved classification.
We use the NLTK library to split the document into individual sentences. Each sentence is then processed through a traditional $N$-layer Transformer Encoder. This results in the start token ($[CLS]$) representing the entire sentence’s information, which encapsulates the semantic features.
As the encoder extracts the content features, the same sentences will be processed using a BERT model pre-trained on PDTB-annotated texts to identify discourse features. This process involves taking two consecutive sentences as input and producing a single discourse relation as output. The first sentence will generate one $[CLS]$ token, and the following sentence will correspond to one discourse relation, referred to as a ‘discourse code’, representing the relationship between itself and its preceding sentence. This module will provide $n$ discourse codes to represent the structure of the document.
After preparing the structural and content features, the code sequences will be input into the decoder. The decoder is similar to a standard Transformer decoder block, with modifications made to cross-attention mechanism. We use $[CLS]$ to generate $Q$, $Code$ to generate $K,V$. The final layer of the decoder will output a probability using a linear network layer and the softmax function to determine the final classification as either MGC or HPC.
6 Implementation Details
We prepared the M4, paraLFQA, and paraWP datasets by randomly splitting each into 80% for training, 10% for validation, and 10% for testing with a fixed random seed. The Plagbench dataset was randomly divided into 60% for training, 20% for validation, and 20% for testing due to the small size of the Plagbench dataset, the standard 10% test split makes it difficult to accurately measure the model’s performance. Since the MGC is generated from HPC, we ensure that the MGC and HPC are split equally to achieve a balanced dataset. Additionally, our MhBART model requires long-form paraphrased texts paired with their original HPC counterparts, with similar meanings between the HPC and MGC. Therefore, we use the M4 dataset only for DTransformer evaluation, specifically utilizing the M4-WikiHow subset, which contains long documents.
We built the RoBERTa-large baseline using the pre-trained model from Hugging Face [^8] https://huggingface.co/openai-community/roberta-large-openai-detector. We adjusted the maximum sentence length from 512 to 1024 to accommodate long-form texts. The model was fine-tuned with a batch size of 2 and a learning rate of 2e-5.
For MhBART and DTransformer, we initialized our model parameters using the bart-large-cnn model [^9] https://github.com/inferless/Facebook-bart-cnn and Kaiming uniform initialization He et al. (2015), respectively. The hyperparameters and optimizer selection are listed in Appendix F.
7 Evaluation
We present the classification accuracy for both our models, alongside the results from the baseline and SOTA academic and commercial approaches, namely RoBERTa Large, GPTZero, Copyleaks, and Winston AI. Given that the dataset is balanced (50% MCG and 50% HPC), accuracy is sufficient as an evaluation metric (a majority class baseline would achieve 50% accuracy). The results are shown in Table 1. Although AUC-ROC is widely used in detection tasks, it is particularly valuable for measuring performance when the cases are imbalanced while our dataset are balanced dataset. We present an AUC-ROC graph for the paraLFQA dataset as an example, demonstrating results consistent with the accuracy metric shown in 5.
Figure 5: AUC-ROC line for paraLFQA dataset of 6 models. It shows that the DTransformer has the highest AUC-ROC score.
The RoBERTa Large detector classified most cases as HPC, resulting in an accuracy around 50%. Subsequently, we discuss the results and analyze the effectiveness of our proposed models.
Table 1: Binary classification accuracy for MhBART and DTransformer compared with baseline and SOTA commercial models on both datasets.
Model paraLFQA paraWP Revised paraWP Plagbench M4-wikihow
GPTZero 0.52 0.95 0 0.69 0.57
Copyleaks 0.50 0.67 0.4 0.59 0.91
Winston ai 0.51 0.50 0.6 0.43 0.93
RoBERTa Large 0.50 0.50 0.2 0.54 0.49
MhBART (BLEU) 0.54 0.78 0.8 0.63 -
MhBART (Cosine) 0.53 0.75 - 0.72 -
DTransformer (PDTB2.0) 0.67 0.99 0.7 0.69 0.95
DTransformer (PDTB3.0) 0.70 0.99 0.7 0.61 0.94
7.1 MhBART analysis
MhBART outperforms RoBERTa Large on both document-level paraphrased datasets as it can be seen in Table 1 in bold score, and performs better on the Plagbench and paraLFQA dataset across all other detectors, achieving 0.72 and 0.54 accuracy, respectively. This indicates that MhBART can learn some implicit features, such as writing styles, in the output text. The output texts tend to have a writing tone that aligns with their related HPC versions. Examples are in Table 2. We hypothesize that the Plagbench dataset may exhibit more distinct writing style features, making it easier for MhBART to learn and thereby achieving the best performance when using cosine similarity as the evaluation metric. For the other datasets, MhBART performs slightly worse than DTransformer, as these datasets contain relatively long texts that DTransformer may be better at taking advantage of the discourse relations. Additionally, BLEU is better suited for capturing format similarity. Beyond, the complexity of these datasets may pose challenges for BERT in encoding text for the cosine similarity calculation.
Table 2: Output text examples. HPC and MGC given as inputs to MhBART and their outputs. For HPC, the output retains more content, whereas for MGC, the output tends to use fewer “ing" forms and has a more human-like writing style.
Original Texts Output Texts
HPC: The wet marble floor pressed on his cheek like a thousand hands slapping his face frozen in time. The wet marble floor pressed on his cheek like a thousand hands slapping his face frozen in time.
MGC: He woke up on the wet marble floor, feeling the discomfort of the cold surface against his cheek. He woke up on the wet marble floor. The floor was cold against his cheek, and he winced at the sensation.
For the paraWP dataset, MhBART may not be the best model, possibly due to the issue with the classification component that it cannot find the threshold of binary classification bound. There might be better classification models to find this threshold, which could provide higher detection accuracy. We provide an additional plot to illustrate the difference in BLEU1 scores between the output of MhBART and the input texts (MGC or HPC). This is shown in Appendix D, suggesting the ineffectiveness of the classification model.
Another possible reason is because paraWP has explicit features that are easily detected by GPTZero. These features include, but are not limited to, the following writing styles: a) using present or past participles as adverbial clauses and using appositives, and b) using metaphors and parallelism in sentences. To test this hypothesis, on a sample of data we removed these writing patterns by reducing the number of adverbial clauses with -ing and -ed forms, appositives, and parallel sentences. This simplifies sentences by using essential logical connectors (such as so, but, because, and) and by deleting repeated elements. We did that on 10 paragraphs (3 are long documents), and it turns out that the GPTZero cannot classify the revised version correctly but MhBART would not be influenced negatively using BLEU-1/4 as a similarity metric. One of the revised example comparisons is shown in Table 6. It shows that the MhBART learns more than simple explicit features which GPTZero might rely on for detection.
We think another factor contributing to the performance of commercial tools is the dataset-specific bias in their outputs. Evidence for this is seen in the lack of a single dominating commercial tool across all of datasets. The exceptional performance of certain tools may be attributed to the specific training samples they were exposed to. For instance, GPTZero might have been trained on a larger number of samples generated by ChatGPT using prompts similar to those used to produce paraWP, but fewer for M4. Consequently, the model’s performance may not necessarily reflect an effective feature extraction strategy, but rather the effectiveness of its training data with specific types of prompts.
7.2 DTransformer analysis
The results presented in the Table 1 demonstrate that the DTransformer outperforms all other models across the paraLFQA, showing its efficacy in extracting features from document-level texts. ParaLFQA is discourse paraphrased so DTransformer could take advantage of discourse inforation while other models fail to pay attantion to. For the Plagbench dataset, it performs slightly worse than MhBART, likely because the dataset lacks a sufficient number of distinct discourse features based on the prompts used for data generation. These performances highlight the significance of incorporating discourse features. The critical of discourse features is further validated through comparative experimentation using PDTB 2.0 and 3.0, where PDTB 3.0 consistently outperforms 2.0. This superiority is attributable to PDTB 3.0’s more comprehensive discourse annotations, which better represent discourse features effectively. Specifically, the best-performing DTransformer model utilizing PDTB 3.0 shows a 4%, 15.5% 1.5% absolute improvement over existing commercial models on the paraWP, paraLFQA, and M4 datasets, respectively.
Our result reveals that DTransformer exhibits high performance on the paraWP dataset, achieving an accuracy of 99%. This indicates that DTransformer possesses a remarkable capability for detecting MGC, even within the constraints of a challenging dataset. Furthermore, the design of the generation prompt plays a crucial role. Specifically, we ensured that the generated MGCs were predominantly distinct from their corresponding HPCs at the discourse feature level, while controlling for other variables such as semantic meaning and word count. This alignment with DTransformer’s capabilities accounts for the exceptionally high accuracy observed.
We also performed a case study, shown in Appendix E to assess the effectiveness of the PDTB.
In addition, we conducted an ablation study focusing on discourse features. To successfully extract document-level semantic meaning as DTransformer does, we trained the original HiBERT model Zhang et al. (2019) on the paraWP dataset. However, the predicted accuracy with the hierarchical architecture model alone is 0.53 and 0.49
for the PDTB3.0 and PDTB2.0 annotated datasets, respectively on paraWP. This finding highlights the importance of incorporating both discourse and semantic features for effective document-level MGC detection. Additionally, it highlights that the accuracy gain for DTransformer is primarily attributable to discourse features.
In light of the models’ high performance, we aim to evaluate the scalability of our models. We conducted five independent experiments for each model, utilizing different random seeds. DTransformer with PDTB 2.0 resulted in a mean test accuracy of 0.67 and a standard deviation of 0.03 on the paraLFQA dataset. The DTransformer with PDTB 3.0 produced a mean test accuracy of 0.70 and a standard deviation of 0.02. These results indicate the high mean accuracy and consistency of both models, demonstrating good scalability. The lower standard deviation of PDTB 3.0 suggests more stable performance across different runs.
8 Conclusions
In this work, we introduced two new datasets aiming to provide insights into the challenges in current work aimed at detecting MGC, particularly for document-level paraphrased machine-generated texts: paraLFQA, emphasizing the discourse difference, and paraWP, a long dataset with highly similar between HPC and MGC. Further, we introduced two architectures, MhBART and DTransformer, outperforming current commercial and non-commercial models. The former incorporates human-written styles, the latter utilizes discourse features. We reported that our models outperform the competition across the datasets paraLFQA, paraWP, Plagbench, and M4. Additionally, we identified patterns in writing styles that affect the performance of existing detectors and show that our detectors are more robust when faced with these patterns. Future work could focus on integrating both models into a end-to-end detection system, enhancing efficiency in detection workflows.
Limitations
A. We only report three sentence patterns, or writing styles, that affect GPTZero’s detection. It could be a future direction to find ways to embed these writing styles into automatic detection methods.
B. We excluded non-latin alphabet languages, such as Tibetan, as the current focus is on certain behaviour of models in a simple testbed without complications. The intention is to avoid complications arising from challenges associated with non-latin alphabet languages. During our study, every text with non-latin characters was removed from the dataset.
C. We eliminated obvious differences between HPC and MGC regarding writing habits. This includes instances of consecutive emotional markers like “!!!!!!” or artificial modifications of words like “su–per!”, used merely to express a pause. We detect if there are invalid markers within the words, then, we removed this sample, as the goal of the study was to specifically focus on learning intrinsic features, such as text structure and writing style.
D. We attempted to embed methods for automatically removing sentence patterns that fool GPTZero, such as using GPT-3.5 and GPT4/4o to rephrase the entire corpus. However, any prompts we tried such as “Rewrite the paragraph to replace all the appositive and participle clauses including ing or ed. Do not change, add or delete content words. Retain other parts of the language expression.” did not remove all the present or past participle clauses. This might be as certain prompts cannot apply to all texts.
E. Due to GPU limitations, we selected a manageable batch size and maximum sentence length for our dataset. As a result, our model has been trained on the target dataset but would benefit from additional training for broader, out-of-domain commercial applications. While commercial models have likely not been exposed to the specific datasets we investigated, they leverage significantly larger datasets in their training processes, though the specifics of this data are not publicly disclosed.
Ethics Statement
Our work has important ethical implications. The advancements in MGC detection not only help to foster a more honest academic environment but also align with the ethical responsibility to maintain integrity in education. As conscious testing for MGC shows initial success, continuing this trend supports academic honesty and fairness for all students. Moreover, effective MGC detection plays a crucial ethical role in combating the spread of fake news and misleading information, which promotes truth and trust in public discourse. In terms of the innovation of music and drawings, detecting MGCs helps preserve the authenticity of artistic expression, preventing forgery and ensuring rightful appreciation to creativity. Through these efforts, we aim to uphold ethical standards across various fields and mitigate the misuse of MGC.
Appendix A Writing prompts data linguistic language features
We investigate basic linguistic features of paraWP: most commonly used words, sentence lengths, lexical diversity, token-type ratio, and further more. The graphs below indicate the top 20 frequently used terms in the HPC (Fig. 6) and MGC (Fig.  7) datasets. It could be seen that the most frequently used words are similar and the number of times the words used is also similar. There is not much difference in these feature.
Figure 6: Top 20 most used word frequencies calculated in the HPC corpus.
Figure 7: Top 20 most used word word frequencies calculated in the MGC corpus. It is used to compare the diversity of words between MGC and HPC.
Additionally, other linguistic features are presented in Tables 3, 4, and 5.
Data Avg. Sent. Length Lexicon Diversity
MGC 486.34 0.498
HPC 559.44 0.474
Table 3: Average Sentence Length and Lexicon Diversity. We calculate the average number of words in one document (Average Sentence Length) and the proportion of different words in total words (Lexicon Diversity). These are two other linguistic features for MGC and HPC linguistic features comparison.
Data Avg. Sent. ing PTCP ed PTCP CC
MGC 28.04 13.57 10.84 16.14
HPC 41.75 11.75 9.93 17.84
Table 4: Average number of sentences, present & past participle and connectives in the test set. We calculate the average number of sentences in one document (Avg. Sent.) and also for present, past participles, and connectives given their total number of characters are similar. These are the comparisons of the sentence patterns between HPC and MGC.
The POS tag ratios provide syntactic information, and the average lexicon diversity is calculated by dividing the number of distinct words used by the total number of words. There are no significant differences in any aspect. Overall, explicit features such as syntactic or content features of MGCs and HPCs are similar. It is challenging to detect MGC or HPC relying only on these.
POS MGC Ratio HPC Ratio
Det 0.1099 0.0993
Noun 0.2424 0.2209
Adj 0.0613 0.0674
Verb 0.2129 0.2235
Adp 0.1156 0.1045
Pron 0.1217 0.1346
Num 0.0051 0.0079
Conj 0.0379 0.0349
Prt 0.0394 0.0356
Adv 0.0531 0.0698
X 0.0007 0.0018
Table 5: POS Tag Ratios for MGC and HPC. We use universial POS tags for word classification and calculate their respective proportions. This serves as an indicator of superficial syntactic features.
Appendix B Universial authorship representation (UAR) representation experiments
The UAR is utilized as implicit features of writing style, proposed as stylistic representation. We calculate the difference score of cosine similarity based on UAR, and the results are quite poor, with accuracy scores of 0.623, 0.574, and 0.518 for paraWP, Plagbench, and paraLFQA, respectively. We visualize the UAR using UMAP, following the methodology proposed in the UAR work Soto et al. (2021), as shown in Fig. 8.
Figure 8: UMAP visualization of UAR representations for four types of texts: HPC (top-left), intermediate output with HPC and MGC as input (top-right), and MGC (bottom-left) in paraLFQA (up) and paraWP (down). Each plot shows approximately 300 texts for each type of texts.
This figure provides insight into the suboptimal performance of MhBART when using UAR. In paraLFQA, there is a lack of distinct separation among the four text types, whereas in paraWP, the intermediate outputs exhibit a higher degree of similarity. This phenomenon is attributed to the limitations of UAR in effectively capturing the entire writing style and its challenges in fine-tuning for our specific task. Future work should focus on developing an automated method to evaluate this implicit feature.
Appendix C Revised MGC example
The comparison can be seen in Table 6. Changes are marked in font in the texts below. Long texts without changes are indicated by . If the AI percentage surpasses 50%, GPTZero should classify the content as MGC. Conversely, if it is 50% or lower, GPTZero should classify it as HPC.
Table 6: Revised MGC example. The font parts indicate all the manual revisions. The entire text is around 450 words, and we show all changes within significant contexts. Parts that are not shown have not been changed.
Original MGC: 100% AI detected by GPTZero Revised MGC: 44% AI detected by GPTZero
He woke up on the wet marble floor, feeling the discomfort of the cold surface against his cheek. Ignoring the rain outside, he struggled to gather his thoughts, but the memories remained veiled by a black mist. Realizing he didn’t know where he was, he searched the room and found a bottle of vodka. Taking a few swigs, he used it as a distraction from his forgotten past. With nothing else to do, he looked at himself in the mirror. His appearance suggested he could be a successful mattress salesman or a lucky gambling bum. His face was unshaven, his stomach round yet firm, and his hair long and unruly. However, his eyes still held a hint of intensity, like a rusted dagger that could still cut. In frustration, he cursed the name "DiCaprio," resenting the fact that despite his efforts to bury himself in alcohol and drugs, he was still associated with the actor. Even after bankruptcy, the royalties from Titanic and other projects kept pouring in, making it impossible for him to escape his past. He threw the bottle at the mirror, but it slipped from his wet hands and only made a thud against the wall. Defeated, he slumped against the floor and noticed the rain entering the room. Frustrated and confused, he wondered what had happened the night before. As he sat there, the pieces of his fragmented memory started to come together. He had been involved in a violent altercation, and the consequences were now staring back at him through the broken window. Fear and regret washed over him as he tried to piece together the events leading up to this moment. The pain in his body matched the pain in his mind, and he knew he couldn’t escape the consequences of his actions. The room, once a place of mystery, now became a haunting reminder of the darkness within him. He slowly stood up, his body aching with every movement. Determined to uncover the truth, he stepped over the broken glass and ventured out into the rain-soaked streets. The journey to unravel the mysteries of that fateful night had just begun, and he knew he had to confront his past, no matter how painful it might be. He woke up on the wet marble floor, and felt the discomfort of the cold surface against his cheek. He ignored the rain outside. He struggled to gather his thoughts, but the memories remained veiled by a black mist. Because he realized he didn’t know where he was, he searched the room and found a bottle of vodka. He took a few swigs, and used it as a distraction from his forgotten past. He looked at himself in the mirror but had nothing else to do. His appearance suggested he could be a successful mattress salesman or a lucky gambling bum. His face was unshaven. His stomach was round yet firm. His hair was long and unruly. However, his eyes still held a hint of intensity, like a rusted dagger that could still cut. He was in frustration, and cursed the name "DiCaprio." This resents that despite his efforts to bury himself in alcohol and drugs, he was still associated with the actor. Even after bankruptcy, the royalties from Titanic and other projects kept pouring in. This made it impossible for him to escape his past. He threw the bottle at the mirror, but it slipped from his wet hands and only made a thud against the wall. He was defeated so he slumped against the floor. He noticed the rain entering the room. He was frustrated and confused. He wondered what had happened the night before. He sat there. The pieces of his fragmented memory started to come together. He had been involved in a violent altercation, so the consequences were now staring back at him through the broken window. Fear and regret washed over him when he tried to piece together the events. The pain in his body matched the pain in his mind, but he knew he couldn’t escape the consequences of his actions. The room is a place of mystery, but now became a haunting reminder of the darkness within him. He slowly stood up. His body aches with every movement. He was determined to uncover the truth, so he stepped over the broken glass and ventured out into the rain-soaked streets. The journey to unravel the mysteries of that fateful night had just begun. He knew he had to confront his past, no matter how painful it might be.
Appendix D BLEU plots for MhBART results
This appendix provides additional plots to show the difference score (BLEU1) between the output of MhBART and input (MGC or HPC texts) shown in Fig. 9
Figure 9: The BLEU score as a difference score for 128 cases in the validation set of WritingPrompts shows the differences in divergence between HPCs and MGCs with the outputs of MhBART.
This evidence suggests that the model learns features from HPCs because the output tends to mimic and resemble HPCs, given that the overall bleu line trend of HPCs is higher than the red line of MGCs, even though there is no clear threshold. There might be better classification models to find this threshold, which could provide higher detection accuracy.
Appendix E Case study for the DTransformer result
We performed a case study, as shown in Table 7, to assess the effectiveness of the PDTB. For this evaluation, we employed the DTransformer with PDTB 3.0 as the representative model. The case study analyzed a paragraph from the paraLFQA dataset, originally labeled as MGC. This paragraph was misclassified by GPTZero and MhBART as HPC, but it was correctly classified as HPC by DTransformer.
Table 7: DTransformer test example: A paragraph from paraLFQA is selected and its corresponding PDTB code prediction is presented. DTransformer made correct classifications while MhBART and GPTZero classified falsely.
Text PDTB code
Basically, they are related to the amount of “advantage" which a lever and a winch give you. [CLS]
A lever or a pulley can make a heavy object move, because they transform a weak but large movement into a small but strong movement. Expansion.Level-of-detail
For example, by operating a come-along, you move a meter of the long lever with your arm, which is transformed into moving the heavy load one centimeter. Expansion.Instantiation
If you do this (divide the distance by a hundred) you multiply the force by a hundred, and so you can handle a weight you could move with 50 kilograms by hand, but with the help of the come-along you can handle a weight of 5,000 kilograms. Expansion.Level-of-detail
When using a ratchet strap to secure a load, the lever you move is much shorter and therefore moves less. Expansion.Level-of-detail
But you have to know that if the end of the lever moves ten centimeters and the ratchet strap moves one centimeter, your force is multiplied by ten. Contingency.Cause
So you convert your strength of 50 kilograms into a strength that can lift 500 kilograms. Contingency.Cause
When the strap is pulled to the point where it feels as if 500 kilograms are pulling on it, your strength of 50 kilograms is no longer enough to tighten it any more. Expansion.Level-of-detail
So, in essence, a come-along has a longer lever to give you a larger advantage. Contingency.Cause
The PDTB codes are intuitively accurate. For instance, the ‘expansion.level-of-detail’ code corresponds to the elaboration on the ‘lever’ in the second sentence, following the ‘lever’ discussed in the previous sentence. Additionally, the ‘contingency.cause’ code is marked by sentences containing conjunctions such as ‘but’ and ‘so’ to indicate cause and effect relationships. The ‘expansion.instantiation’ code is identified by the presence of the phrase ‘for example.’ These examples demonstrate that PDTB coding aligns well with human judgment, providing evidence for the effectiveness of discourse features in detecting MGC. Conversely, we hypothesize that MhBART and GPTZero produced incorrect classifications due to their relative emphasis on writing style and other semantic features rather than on discourse features.
Appendix F Hyperparameter and opmimizer for our models
Table 8: Hyperparameters, tokenizer and optimiser for MhBART training.
Hyperparameter MhBART
max-sentence-length 1024
batch 1
learning rate 2e-5
epochs 10
early stop epochs 3
optimizer AdamW
tokenizer BPE
Hyperparameter Value
embedding dim 512
feedforward embedding dim 2048
encoder layer 6
decoder layer 6
encoder attention heads 8
decoder attention heads 8
dropout 0.1
activation function sigmoid
max-sentence-length 128
max-paragraph-length 128
batch 4/16
learning rate 2e-6
epochs 20
early stop epochs 5
optimizer Adam
tokenizer white-space
Table 9: Hyperparameters, tokenizer and optimizer for DTransformer.
The batch sizes differ between the two datasets of DTransformer training is due to GPU memory limitations. We use a batch size of 4 for paraWP and 16 for paraLFQA, reflecting the varying dataset sizes, with paraLFQA being slightly shorter in length. The device we use is a Tesla V100-PCIE-32GB. Also, we could only use batch size 1 for the MhBART experiments due to the GPU memory limitations. This constraint likely contributes to the suboptimal performance of our MhBART model, as batch size plays a crucial role in model training.